# AI-Driven Mobility Management for High-Speed Railway Communications: Compressed Measurements and Proactive Handover

### v2

@@ III-B AI-Based Proactive Handover Scheme
Existing AI/ML-assisted cell handover methods typically use the RSRP of all beams across cells to predict the future RSRP for each beam or to identify the optimal base station index for upcoming handovers [15, 3]. These methods primarily focus on the improvements in prediction accuracy and spectral efficiency provided by AI models, with limited attention to system-level metrics, such as RLF rates and PP rates. This gap may overlook the potential impact of model design on system reliability and service continuity. Additionally, existing algorithms do not sufficiently address the inherent limitations of the event A3-based handover mechanism. For instance, event A3-based handover is triggered and executed based on historical measurement reports, making it a passively reactive approach. The system initiates the handover only after the UE detects communication quality degradation, where the communication quality has already suffered irreversible loss.
To maintain a stable communication quality for HSR communications, we provide a novel AI-assisted cell handover scheme. We use an AI model to predict the L3-RSRP of all cells for future state 1 based on historical L1-RSRP to aid the source BS in handover decisions. Using this method, we predict the future communication quality variation trends of adjacent cells and proactively switch cells based on these predictions. This approach effectively mitigates the communication quality degradation and reduces handover time from the original states (state 1, state 2, state 3) to the existing states (state 2, state 3). The AI-based cell handover process is illustrated in Fig. 5, where option 1 and option 2 represent different deployment methods of AI models. In option 1, the AI-based prediction model is deployed on the UE side, with the prediction report uploaded to the source BS to assist in the handover decision. In option 2, the AI-based prediction model is deployed on the network side, where the network performs temporal prediction and makes the handover decision based on historical measurement reports. The primary difference between the two options lies in the model’s deployment mode, while the model’s inputs, outputs, and architecture remain the same.
@@
Based on the handover scheme we propose, we identify two approaches to reduce measurement overhead in cell-level mobility management. For the first approach, all beams from the partially selected cells are used to predict the L3-RSRP for all cells. For the second approach, the partially selected beams of each cell are used to predict the L3-RSRP for all cells. To be noticed, the model input is obtained through equal-interval sampling rather than the compressed multi-beam measurement method proposed in Section II to independently evaluate the performance gains of the AI-assisted handover scheme. We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances. We employ CNN-based and LSTM-based models for RSRP prediction. The LSTM-based model comprises four LSTM layers and two FC layers. To enhance the performance of the LSTM network, we adopt four identical networks to predict L3-RSRP for four future instances separately. For the CNN-based model, a set of regular 2D convolution layers with 12, 24 and 32 kernels are employed before feeding it into two FC layers. To avoid overfitting during the training process, max pooling layers are incorporated after each CNN layer.
@@
The channel used in our experiments is generated by QuaDRiGa, with the UMa scenario and a center frequency of 30 GHz, consistent with the beam-level experiment setup. We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees and interference is generated by 21 adjacent cells, modeled as random interference. We generate four datasets corresponding to various UE speeds of 60 km/h, 120 km/h, 350 km/h, and 500 km/h. Each dataset comprises 50,000 samples, which are split into training, validation, and test sets in a 3:1:1 ratio. For the proposed two approaches with equidistant sampling, L1-RSRP of 50% beams in each cell or all beams in 50% cells are selected as input to the AI model. The batch size is 256 and the maximum epoch is 800. The learning rate dynamically changes from $10^{-3}$ to $10^{-8}$, utilizing the adaptive moment estimation optimizer. The RSRP difference is one evaluation metric of prediction accuracy for cell-level mobility management, which reflects the disparity between the predicted L3-RSRP and the actual L3-RSRP of all cells. In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system. The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance. Table II shows the significantly reduced RLF rates in AI-assisted cell handover processes. Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications. Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model, while achieving comparable performance.
@@ IV-A Dynamic Interference Management
Due to the high mobility in HSR communications, the randomness of inter-cell interference constitutes a crucial challenge. Specifically, mmWave beams from neighboring cells directed at local terminals can cause severe inter-cell interference, while frequent cell handovers and beam switching exacerbate the randomness of this interference. Therefore, agile interference management is crucial to ensure reliable service in HSR communications. As a potential solution, graph neural networks (GNNs) can intrinsically match the topology of wireless networks and can be used to effectively handle dynamic interference management.
@@ IV-B Multi-Criteria Joint Decision-Making Mechanism for Mobility Management
In HSR environments, the rapid movement necessitates more adaptable mobility management strategies to effectively handle frequent environmental changes. Relying solely on RSRP variations is insufficient for ensuring communication stability during handover decisions. Therefore, a multi-criterion decision-making mechanism becomes imperative. This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs to the AI models to obtain better handover decisions. This would greatly improve the accuracy and reliability, ensuring seamless communication for high-speed trains.
@@ IV-C Model Deployment and Compression
In practical scenarios, the AI-assisted mobility management methods require high energy dissipation due to the high computing complexity, so it is more meaningful to deploy these AI models on the network side. In this case, the UE reports measurement data to the source BS, which feeds the measurements into the AI/ML models for mobility management. Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring during the reporting of these measurements, affect the performance of AI models. Then, it would be beneficial to deploy AI models on the UE side. We can consider employing advanced AI models, such as lightweight AI models or transformer-based architectures, to enhance the computational efficiency and reduce the hardware requirements for the UEs.
@@ IV-D Model Generalization Design
AI/ML models designed for specific scenarios would usually have poor generalization performance. Altering the settings in communication scenarios, e.g., BS height, channel model, cell size and UE speed, might lead to a significant deterioration of the model performance. Therefore, one challenge for AI-based mobility management for HSR communications is the design of AI models with good scenario generalization properties. Continuous learning is particularly beneficial in dynamic environments, enabling models to adapt to shifting channel conditions and movement patterns in real time. However, enhancing the model generalization through continuous learning incurs significant deployment overhead. Balancing deployment overhead and generalization remains a critical research in HSR communication.
@@ V Conclusion
In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications of AI models. For beam-level mobility management in HSR communications, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead. For cell-level mobility management in HSR communications, we unveiled the significant gain of AI-based methods for cell handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates. At last, the opportunities and challenges on AI-based mobility management were briefly discussed for further studies.

### v3

@@ III-B AI-Based Proactive Handover Scheme
Existing AI/ML-assisted cell handover methods typically use the RSRP of all beams across cells to predict the future RSRP for each beam or to identify the optimal base station index for upcoming handovers [15, 3]. These methods primarily focus on the improvements in prediction accuracy and spectral efficiency provided by AI models, with limited attention to system-level metrics, such as RLF rates and PP rates. This gap may overlook the potential impact of model design on system reliability and service continuity. Additionally, existing algorithms fail to address the limitations of the event A3-based handover mechanism, which is a passively reactive approach triggered only after the UE detects irreversible communication quality degradation.
To maintain a stable communication quality for HSR communications, we provide a novel AI-assisted cell handover scheme. We use an AI model to predict the L3-RSRP of all cells for future state 1 based on historical L1-RSRP to aid the source BS in handover decisions. Using this method, we predict the future communication quality variation trends of adjacent cells and proactively switch cells based on these predictions. This approach effectively mitigates the communication quality degradation and reduces handover time from the original states (state 1, state 2, state 3) to the existing states (state 2, state 3). The AI-based cell handover process is illustrated in Fig. 5, where option 1 and option 2 represent different deployment methods of AI models. In option 1, the AI-based prediction model is deployed on the UE side, with the prediction report uploaded to the source BS to assist in the handover decision. In option 2, the AI-based prediction model is deployed on the network side, where the network performs temporal prediction and makes the handover decision based on historical measurement reports. The primary difference between the two options lies in the model’s deployment mode, while the model’s inputs, outputs, and architecture remain the same.In HSR scenarios, option 1 is more suitable as it directly uploads the predicted results instead of raw measurement data, reducing the transmission overhead and saving time-frequency resources.
@@
Based on the proposed handover scheme, we explore two approaches to reduce measurement overhead. One uses all beams from partially selected cells, while the other uses partially selected beams of each cell to predict the L3-RSRP for all cells. The model input is obtained through equal-interval sampling rather than the compressed multi-beam measurement proposed in Section II to independently evaluate the performance gains of the AI-assisted handover scheme. We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances. Meanwhile, we employ CNN-based and LSTM-based models for RSRP prediction. The LSTM model captures temporal patterns in handover-related data and consists of four LSTM layers and two FC layers, with four parallel networks predicting L3-RSRP at different future instances. The CNN model, offering lower complexity and faster inference, includes 2D convolution layers with 12, 24, and 32 kernels, followed by two FC layers. While this study adopts classical AI architectures such as LSTM, CNN, and Conv-LSTM to validate the effectiveness of AI-driven mobility management, these models can be replaced in the future by more advanced architectures—such as large language models—to further enhance the performance of AI-driven beam-level and cell-level mobility management.
@@
The L1-RSRP and L3-RSRP in our experiments are generated using a system-level simulation platform introduced in Section II.B. We have uploaded two datasets to GitHub with detailed descriptions and scenario configurations. [^6] https://github.com/wc253/dataset_mobility_management Beam-level and cell-level setups differ mainly in how they handle interference. Beam-level experiments focus on intra-cell beam switching and disregard inter-cell interference. In contrast, cell-level experiments involve inter-cell handover and account for inter-cell interference by modeling it stochastically in our setup. Each dataset comprises 50,000 samples, which are split into training, validation, and test sets in a 3:1:1 ratio. For the proposed two approaches with equidistant sampling, L1-RSRP of 50% beams in each cell or all beams in 50% cells are selected as input to the AI model. The batch size is 256 and the maximum epoch is 800. The learning rate dynamically changes from $10^{-3}$ to $10^{-8}$, utilizing the adaptive moment estimation optimizer. The RSRP difference is one evaluation metric of prediction accuracy for cell-level mobility management, which reflects the disparity between the predicted L3-RSRP and the actual L3-RSRP of all cells. In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system. The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance. Table II shows the significantly reduced RLF rates in AI-assisted cell handover processes. Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications. Furthermore, as shown in Table II, the number of flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model, while achieving comparable performance.
@@ IV-A Dynamic Interference Management
Due to the high mobility in HSR communications, the randomness of inter-cell interference constitutes a crucial challenge. Specifically, mmWave beams from neighboring cells directed at local terminals can cause severe inter-cell interference, while frequent cell handovers and beam switching exacerbate the randomness of this interference. Thus, agile interference management is essential for reliable service. As a potential solution, graph neural networks (GNNs) can intrinsically match the topology of wireless networks and can be used to effectively handle dynamic interference management.
@@ IV-B Multi-Criteria Joint Decision-Making Mechanism for Mobility Management
In HSR environments, rapid movement demands adaptive mobility management strategies to handle frequent environmental changes. Relying solely on RSRP variations is inadequate for stable handover decisions. A multi-criterion decision-making mechanism is essential, incorporating factors such as signal interference, network load, terminal velocity, and movement direction alongside RSRP. Integrating these inputs into AI models enhances handover accuracy and reliability, ensuring seamless communication for high-speed trains. Future research could further improve the multi-criteria joint decision mechanism by exploring advanced optimization techniques, such as reinforcement learning-based adaptive weighting of criteria, or leveraging federated learning to enhance model generalization across different HSR environments.
@@ IV-C Model Deployment and Compression
AI-assisted mobility management in practical scenarios demands high energy consumption due to its computational complexity, making network-side deployment more viable. In this setup, the UE reports measurement data to the source BS, which processes it using AI/ML models for mobility management. However, the transmission introduces additional latency, and inevitable reporting errors can degrade the AI model‘s performance. Deploying AI models on the UE side could mitigate these issues. Leveraging advanced AI models, such as lightweight architectures or transformer-based designs, can enhance computational efficiency while reducing UE hardware requirements.
@@ IV-D Model Generalization Design
AI/ML models designed for specific scenarios would experience performance degradation when applied to different railroad environments due to variations in factors such as BS height, channel characteristics, and railway layouts. To mitigate this issue, the most effective approach is to collect new datasets from the target environment and either retrain the model or apply transfer learning to fine-tune it. This enables the model to better adapt to the new conditions while minimizing performance loss. Furthermore, continuous learning is crucial in dynamic environments, allowing models to adapt to changing channel conditions and movement patterns in real time. However, improving model generalization through continuous learning introduces significant deployment overhead. Balancing these factors remains a key challenge in HSR communication.
IV-E Model Explainability
Explainability is essential to ensure model trustworthiness, safety, and effectiveness. However, achieving explainability in HSR communications presents several challenges. First, HSR communications demand high accuracy in decision-making. More complex models usually provide better performance but tend to be more opaque and difficult to interpret. Striking a balance between performance and explainability remains a key challenge. Moreover, complex interpretability methods can increase computational overhead, potentially affecting the communication system’s response time. Therefore, finding a balance between explainability and real-time performance is a critical research direction in HSR communications.
@@ V Conclusion
This article explored AI-based beam-level and cell-level mobility management in HSR communications. Firstly, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead. Secondly, we unveiled the significant gain of AI-based methods for proactive handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates. Lastly, the opportunities and challenges on AI-based mobility management were briefly discussed for further studies.